OpenStudy (music0203):

solve the equation by completing the square. round to the nearest hundredth if necessary. x^2+2x=8

OpenStudy (music0203):

@Michele_Laino @kidrah69

OpenStudy (lυἶცἶ0210):

Well you have \(\large x^2+2x= 8\) Now take b (2) and: \(\large (\frac{b}{2})^2\) = \(\large (\frac{2}{2})^2 \) =\(1^2\) =1 So: \(\Large x^2+2x\color{red}{+1}=8\color{red}{+1} \) Now add 1 to both sides to make it even, now condense: \(\Large (x+1)^2=9 \) Subtract and set it equal to 0: \(\Large (x+1)^2-9 =0\) and that should be it If you're solving then take the square root on both sides: \(\Large \sqrt{(x+1)^2} =\sqrt{9} \) \(\Large (x+1)=\pm 3 \) Now solve for x: \(\Large x=\pm3-1 \) :)